# TableMind: An Autonomous Programmatic Agent for Tool-Augmented Table Reasoning

### v2

@@ 5.4. Comparison of Model Variants
5.5. Analysis of Tool Usage
5.5.1. Tool calls ratio.
The evolution of the agent’s policy for invoking tools is shown in Figure 5 (left). The learning trajectory exhibits a clear and continuous progression with distinct phases. Initially, during the first 40 steps, the agent undergoes an exploration phase, where its commitment to using tools is tentative and fluctuates as it balances direct textual generation with tool interaction, occasionally reverting to text-only outputs when tool use does not appear immediately beneficial. This is followed by a period of rapid discovery and exploitation, spanning steps 41 to 70, marked by a sharp inflection in the curve. During this stage, the agent decisively associates tool invocation with high rewards, consistently reinforcing a policy that favors code generation and beginning to integrate tool outputs more effectively into its reasoning process. Finally, during steps 71 to 120, the agent’s strategy converges, with the tool-use ratio stabilizing near 100%, indicating that the model has learned a near-deterministic policy that treats tool use as the optimal approach for the vast majority of tasks in this domain and applies it with high confidence.
5.5.2. Pass ratio.
The Pass Ratio, illustrated in Figure 5 (right), quantifies the reliability and correctness of the agent’s generated code, measured as the proportion of executions producing the expected results without errors. The curve shows a clear and sustained improvement in code-writing proficiency over the course of RFT training, indicating progressive refinement of the agent’s coding strategy. The model begins with a baseline pass ratio of approximately 60%, reflecting the syntactic and semantic code generation capabilities developed during the SFT stage. During the RFT phase, the pass ratio steadily rises to nearly 90%, with execution-based feedback guiding the agent to reinforce correct coding patterns, eliminate common sources of runtime errors, and align code more closely with task-specific requirements. These results suggest that the combination of SFT and RFT produces an agent that not only decides when to invoke a tool but also executes the invocation reliably and consistently across different problem settings.
@@ 5.6. Hyperparameter Sensitivity Study
We conducted a targeted study to identify the optimal settings for key hyperparameters and evaluate how sensitive our model is to their variations. This study focused on two core hyperparameters: the maximum number of interaction turns (Max Turns) and the generation Temperature. The results are summarized in Table 3. The Max Turns parameter defines the maximum number of thought-action cycles the agent can perform to solve a problem. The results indicate that: Model performance significantly improves across all three datasets when Max Turns is increased from 1 to 3. This suggests that a single turn is insufficient for solving the complex, multi-step reasoning required by these tasks. However, when Max Turns is further increased from 3 to 5, a slight degradation in performance is observed. This is likely because excessive turns can increase the risk of error propagation or the generation of redundant, unhelpful steps, which can interfere with the final outcome. Based on these findings, we identify 3 as the optimal value for Max Turns. The Temperature parameter controls the randomness of the model’s generated outputs. A lower temperature leads to more deterministic responses, while a higher temperature encourages diversity. The results in the lower half of Table show that:Within the tested range, model performance is positively correlated with Temperature. The peak performance is achieved at 1.0.
@@ 5.7. Case Studies
We present a detailed case study in Figure 6 that reveals the inner workings of TableMind on a representative WikiTQ problem. This process clearly demonstrates the model’s advanced reasoning capabilities, centered on its distinctive plan-action-reflect cycle, which sets it apart from more basic, linear approaches. The model begins by formulating a simple initial plan, successfully extracting two time strings from the table. Crucially, instead of proceeding with a naïve computation that would yield an incorrect result, the model enters a reflection state. In this stage, it identifies that a simple subtraction is infeasible because the time data is stored as textual strings rather than numerical values. This self-reflection prevents premature errors and guides the model toward a more robust solution strategy. Armed with this insight, the model devises a refined plan: it first converts the textual strings into a proper time format suitable for arithmetic operations, and then computes the precise difference. To implement this plan, the model generates executable code using Python’s datetime module. The code runs flawlessly, producing the correct numerical output of 192 seconds. Finally, the model synthesizes this raw numerical result into a well-formatted answer, ready to be returned. This example vividly illustrates that TableMind is capable of decomposing complex queries, reflecting on intermediate steps, self-correcting its reasoning strategy, and strategically leveraging external tools to solve table-based problems. This case epitomizes the powerful synergy of self-reflection, precise tool invocation, and proactive error correction, which is fundamental to the model’s advanced reasoning capabilities.
@@ 6. Conclusion
In this paper, we proposed TableMind, a novel programmatic agent for tool-augmented table reasoning that achieves autonomous and adaptive problem-solving. TableMind first decomposes complex problems, then iteratively generates code, executes it using a code interpreter, and reflects on the feedback to dynamically plan its next actions. Optimized through a two-stage framework combining Supervised Fine-tuning with Reinforcement Learning algorithm, TableMind produces reliable and effective solutions, demonstrating a high degree of adaptability akin to a human analyst capable of debugging and refining their approach. Extensive experiments on multiple table reasoning benchmarks demonstrated that TableMind achieves state-of-the-art performance, validating its effectiveness and robustness. This study targeted the development of a more autonomous paradigm for tool-integrated reasoning, balancing foundational knowledge from supervised learning with advanced strategic discovery from reinforcement learning. While TableMind shows strong performance on single-table tasks, extending its capabilities to handle multi-table joins, a wider array of data science tools, and improving the sample efficiency of its RFT training remain exciting directions for future exploration.

### v3

@@ 5.4. Comparison of Model Variants
5.5. Analysis of Tool-use
5.5.1. Tool Calls Ratio Analysis.
The evolution of the agent’s policy for invoking tools is shown in Figure 5 (left). The learning trajectory exhibits a clear and continuous progression with distinct phases. Initially, during the first 40 steps, the agent undergoes an exploration phase, where its commitment to using tools is tentative and fluctuates as it balances direct textual generation with tool interaction, occasionally reverting to text-only outputs when tool use does not appear immediately beneficial. This is followed by a period of rapid discovery and exploitation, spanning steps 41 to 70, marked by a sharp inflection in the curve. Finally, during steps 71 to 120, the agent’s strategy converges, with the tool-use ratio stabilizing near 100%, indicating that the model has learned a policy that treats tool use as the optimal approach for the vast majority of tasks in this domain and applies it with high confidence.
5.5.2. Pass Ratio Analysis.
The pass ratio, illustrated in Figure 5 (right), quantifies the reliability of the agent’s generated code, measured as the proportion of executions producing the expected results without errors. The curve shows a clear and sustained improvement in code-writing proficiency over the course of RFT training, indicating progressive refinement of the agent’s coding strategy. The model begins with a baseline pass ratio of approximately 60%, reflecting the syntactic and semantic code generation capabilities developed during the SFT stage. During the RFT phase, the pass ratio steadily rises to nearly 90%, with execution-based feedback guiding the agent to reinforce correct coding patterns, eliminate common sources of runtime errors, and align code more closely with task-specific requirements. These results suggest that the combination of SFT and RFT produces an agent that not only decides when to invoke a tool but also executes the invocation reliably and consistently across different problem settings.
@@ 5.6. Hyperparameter Sensitivity Study
We conduct a targeted study to identify the optimal settings for key hyperparameters and evaluate how sensitive our model is to their variations. This study focused on two core hyperparameters, the maximum number of interaction turns and the generation temperature. The results are summarized in Table 4. The max turns parameter defines the maximum number of thought-action loops the agent can perform to solve a problem. The results indicate that: model performance significantly improves across all three datasets when max turns is increased from 1 to 3. This suggests that a single turn is insufficient for solving the complex, multi-step reasoning required by these tasks. However, when max turns is further increased from 3 to 5, a slight degradation in performance is observed. This is likely because excessive turns can increase the risk of error propagation or the generation of redundant, unhelpful steps, which can interfere with the final outcome. Based on these findings, we identify 3 as the optimal value for max turns. The temperature parameter controls the randomness of the model’s generated outputs. A lower temperature leads to more deterministic responses, while a higher temperature encourages diversity. The results in the lower half of the table show that, within the tested range, model performance is positively correlated with temperature.
@@ 5.7. Case Studies
We present a detailed case study in Figure 6 that reveals the inner workings of TableMind on a representative WikiTQ problem. The model begins by formulating a simple initial plan, successfully extracting two time strings from the table. Crucially, instead of proceeding with a naïve computation that would yield an incorrect result, the model enters a reflection state. In this stage, it identifies that a simple subtraction is infeasible because the time data is stored as textual strings rather than numerical values. Armed with this insight, the model devises a refined plan: it first converts the textual strings into a proper time format suitable for arithmetic operations, and then computes the precise difference. To implement this plan, the model generates executable code using python’s datetime module. The code runs flawlessly, producing the correct numerical output of 192 seconds. Finally, the model synthesizes this raw numerical result into a well-formatted answer, ready to be returned. This example vividly illustrates that TableMind is capable of decomposing complex queries, reflecting on intermediate steps, self-correcting its reasoning strategy, and strategically leveraging external tools to solve table-based problems.
@@ 6. Conclusion
In this paper, we proposed TableMind, an autonomous programmatic agent for table reasoning. TableMind enabled multi-turn reasoning and precise numerical computation through a two-stage training strategy combining supervised fine-tuning and reinforcement learning. This design allowed the model to generate and execute programs, verify intermediate results, and iteratively refine its reasoning. Experiments across diverse benchmarks showed consistent improvements over existing baselines, demonstrating the advantages of training lightweight autonomous agents over training-free workflow-based approaches. While this work focused on single-table reasoning, extending TableMind to multi-table settings, richer tools, and more sample-efficient reinforcement learning remained promising directions for future research.
Acknowledgments
This research was supported by grants from the National Natural Science Foundation of China (No. 62502486, 62337001), the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), USTC ResearchFunds of the DoubleFirst-Class Initiative (No. YD2150002501), the Fundamental Research Funds for the Central Universities of China (No. WK2150110032).